Please answer fastSaturated liquid at a pressure 200 bar is contained in a piston-cylinder system at a volume of V 0.01 m3. The system undergoe
Answer #1

given Pi= 200bar V = 0.01 m3 Pi= 20 obce z Vi= 0.002038 molkg 3= 0 J 4L=44= 1785-8 KJ 1 Kg m= ï = 0.002038 m= 4,91kgH! 1-2 constant volume process T2=200°C 2112=862.6 kJ/kg Vi = V2-0.002038 malkg JP2= Psat = 1554-9 kPa Uz = Vj tagg 0.00 2038